• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yoomin

    Log entry time 08:01:16 on January26,2009

    Entry number 257833

    keyword=Owl Shift Summary

    Owl Shift Summary - 26 Jan 2009

    Workers: Yoomin Oh, Khem Chirapatpimol

    Everything has been going well but the Compton CODA crashed and
    not recovered.

    ---------------------------------------------------------------
    00:00 Shift start with run 5582(30716)
    00:25 Call from MCC: they have a problem on a cavity-- no beam for
    about 10 minutes. Stopped 5582.
    00:45 Beam back, BPM looks OK, Start run 5583
    01:05 Call from MCC: He cryo level problem-- no beam, stopped 5583
    01:25 Beam back, Start 5584 with HAPPEx 30717
    03:00 Compton CODA crahsed, tried to restart but failed.
    05:10 run 5587 with HAPPEx 30718 start.
    05:50 MCC called they would do some test
    so the beam would be off for ~15 minutes
    after our current production run end.
    06:30 As run 5587 ended, moved the target to empty position and
    now MCC is tuning the beam.
    06:40 MCC finished their tuning, target back to production cell,
    Start 5588 with 2uA, seems fine, ramping up back to 14uA.
    08:00 5588 and 30718 are running.
    ---------------------------------------------------------------
    Production Run Summary
    Run# Time Charge(C) Comment
    5582 23:47-00:24 0.01978
    5583 00:46-01:05 0.01102
    5584 01:25-02:35 0.05498
    5585 02:35-03:52 0.05472
    5586 03:53-05:09 0.05463
    5587 05:12-06:30 0.0547
    5588 06:43- - Good Production Run
    --------------------
    Total Charge 0.24983